### v3

# MiNER: A Two-Stage Pipeline for Metadata Extraction from Municipal Meeting Minutes

Abstract
Municipal meeting minutes are official documents of local governance, exhibiting heterogeneous formats and writing styles. Effective information retrieval (IR) requires identifying metadata such as meeting number, date, location, participants, and start/end times, elements that are rarely standardized or easy to extract automatically. Existing named entity recognition (NER) models are ill-suited to this task, as they are not adapted to such domain-specific categories. In this paper, we propose a two-stage pipeline for metadata extraction from municipal minutes. First, a question answering (QA) model identifies the opening and closing text segments containing metadata. Transformer-based models (BERTimbau and XLM-RoBERTa with and without a CRF layer) are then applied for fine-grained entity extraction and enhanced through deslexicalization. To evaluate our proposed pipeline, we benchmark both open-weight (Phi) and closed-weight (Gemini) LLMs, assessing predictive performance, inference cost, and carbon footprint. Our results demonstrate strong in-domain performance, better than larger general-purpose LLMs. However, cross-municipality evaluation reveals reduced generalization reflecting the variability and linguistic complexity of municipal records. This work establishes the first benchmark for metadata extraction from municipal meeting minutes, providing a solid foundation for future research in this domain.
1 Introduction
Municipal meeting minutes are important documents of local governance, providing official records of decisions, participants, and voting outcomes. Beyond their legal, administrative, and civic importance, these documents are increasingly used for research [5, 12]. Effective information retrieval (IR) from these records requires extracting structured metadata, including meeting number, type of meeting, date, location, participants, and the session times, which are consistently present in parliamentary datasets[7, 11] and meeting minutes. While pre-trained NER models perform well on general-purpose categories (e.g, person, organization, location), they fail to capture the fine-grained, domain-specific metadata typical of these records. Dedicated models for this domain remain, to the best of our knowledge, nonexistent. Therefore, we propose a two-stage pipeline for automatic metadata extraction. First, a question answering (QA) model identifies the opening and closing segments of each document, reducing noise and focusing subsequent processing. Second, transformer-based models perform fine-grained entity recognition, enhanced via deslexicalization to improve generalization across municipalities. We evaluate our approach on a new dataset of 120 Portuguese municipal minutes, comprising a total of 180 metadata segments, with an English automatic translation for cross-lingual testing. Comparisons with open (Phi [1]) and closed (Gemini [23]) weights LLMs regarding predictive performance, inference time and estimated carbon footprint show that specialized transformer models, outperform general-purpose LLMs in every aspects, while revealing the challenges of cross-municipality generalization. The main contribution are 1) a two-stage pipeline, combining QA-based segmentation with transformer-based NER for fine-grained entity recognition and 2) the release of the corresponding dataset [^1] https://github.com/LIAAD/MiNER, and fine-tuned models [^2] https://huggingface.co/collections/liaad/citilink, establishing a benchmark for metadata extraction from municipal minutes.
2 Related Work
Metadata extraction has been explored across various domains, being crucial for indexing, retrieval, and structured analysis in complex documents. Resources such as IsraParlTweet [15] and Watanabe et al. [24] demonstrate the utility of metadata in parliamentary and scholarly contexts, while also highlighting the limitations of LLMs for automatic metadata extraction. However, most approaches assume metadata sections are already identifiable, overlooking the challenge of locating them automatically. This problem can instead be framed as a boundary detection task, for which QA-based formulations  [17, 16] have proven effective in isolating relevant spans and identifying the start and end points of text segments. Following boundary identification, NER can be applied to extract fine-grained entities such as dates, locations, participants, and other administrative details [22, 19]. Our work differs from prior research by focusing on municipal meeting minutes, a domain that exhibits greater stylistic variability and finer-grained metadata categories than parliamentary or scholarly texts. Furthermore, unlike previous approaches that assume metadata is pre-segmented, we explicitly identify metadata segments before entity extraction. This combination of domain-specific segmentation and fine-grained NER makes our approach distinct within the broader landscape of metadata extraction.
3 MiNER: Minutes Named Entity Recognizer Pipeline
Our pipeline has two main phases, transforming a raw document $D$ into a set of labeled metadata entities extracted from the specific document region. Formally, let $D=\{s_{1},s_{2},\dots,s_{n}\}$ be an ordered sequence of sentences, and let $\mathcal{M}=\{m_{1},\dots,m_{K}\}$ denote the set of metadata categories of interest (e.g., meeting number and type, date, location, participants, session times). We decompose the task into two stages as follows.
Stage 1 - Metadata Boundary Detection (MBD). We start by identifying the opening and closing sections of the minutes where metadata is found. This is approached using extractive QA. For each document $D$, we employ two prompts $p_{t}$, one for each segment $t\in\{\text{opening},\text{closing}\}$, which output the positions $\hat{a}_{t}$ and $\hat{b}_{t}$ of the last and the first sentences of the opening and closing segment, respectively. The identified segments are then defined as $S_{t}=\{s_{\hat{a}_{t}},\dots,s_{\hat{b}_{t}}\},\quad t\in\{\text{opening},\text{closing}\},$ and the concatenation $R=S_{\text{opening}}\cup S_{\text{closing}}$ is the reduced text region passed to Stage 2. Formally, the Stage 1 model is defined as
$f_{\text{BD}}(D,p_{t};\theta_{\text{QA}})\rightarrow R,$
where $\theta_{\text{QA}}$ is the learned parameters of the QA model. The QA approach follows a SQuAD v2-style setup [16, 17]. The model outputs either a text span $(\hat{a}_{t},\hat{b}_{t})$ or a null prediction, corresponding to the empty segment.
Stage 2 - Metadata Entity Recognition (MER). Given the reduced text $R$, the goal is to identify and classify all metadata mentions. Formally,
$f_{\text{NER}}(R;\theta_{\text{NER}})\rightarrow\mathcal{E},$
where $\mathcal{E}=\{e_{1},\dots,e_{R}\}$ is the set of extracted entities, each defined by its type $\tau\in\mathcal{M}$ and a consecutive token span $(u,v)$ within $R$, and $\theta_{\text{NER}}$ represents the learned parameters of the NER model. The model is trained using standard token-level classification objectives (BIO or CRF), restricted to $R$ to reduce noise and improve precision. To promote cross-municipality generalization, a deslexicalization function $\phi(\cdot)$ is applied prior to training [20]. In this context, participants and location names are replaced with synthetic values (60% probability) using Faker library [8] and dates and times are randomly varied in format or content (30% probability). Municipality mentions are also replaced with a placeholder token (@MUNICIPIO). The pipeline is composed of the two stages:
$\mathcal{F}(D;\theta_{\text{QA}},\theta_{\text{NER}})=f_{\text{NER}}\big(f_{\text{BD}}(D;\theta_{\text{QA}});\theta_{\text{NER}}\big)\rightarrow\mathcal{E},$
4 Experimental Setup
To evaluate our pipeline, we conducted experiments covering both stages.
Evaluation Metrics. For MBD, we report Exact Match (EM), which gives the proportion of predictions that exactly match the ground-truth span, and F1-score, which measures token-level overlap between predicted and ground truth spans [16, 17]. For MER, we report precision, recall, and F1-score. In experiments involving LLMs, we also measure inference cost and estimated carbon footprint, using Code Carbon library [4]. In addition, cross-municipality generalization was evaluated using a leave-one-out setup, training on five municipalities and testing on the remaining one. We also simulated real deployment with scarce data adding one annotated minute at a time to assess the number of annotated documents required to achieve stable performance.
Dataset. We make use of the newly introduced CitiLink dataset [2], which consists of 120 Portuguese municipal meeting minutes from six municipalities, manually annotated with multiple layers of information [3], with the annotation guidelines publicly available in the repository. In this work, we focus exclusively on the metadata annotations, consisting of 32,364 metadata segments. Annotated entities cover meeting number (120), date (120), location (100), start time (120), end time (100), meeting type (101), and participants (president (119) and councilors (966), including presence status), totaling eight distinct categories. Personal information was anonymized (e.g., replaced with “***”) to ensure privacy and compliance with data handling regulations. For cross-lingual evaluation, an English version was automatically translated using Azure Translator, and annotations were aligned with LinguAligner [14]. The test set in English was further reviewed and corrected by human annotators. The dataset was subsequently adapted for the two stages of our pipeline. For MBD, each document was converted into SQuAD v2-style format, where the full text serves as context, and each prompt (opening or closing segment) is paired with the corresponding answer span. For MER, only the opening and closing segments from Stage 1 were considered. Due to space limitations, additional details are described in the repository.
Models and Baselines. For MBD, we used deepset/xlm-roberta-large-squad2 [6]. Training ran for 3 epochs (lr = 3e-5, batch = 8, weight decay = 0.01, max length = 512, stride = 128). We compared against unsupervised baselines: BM25, [18] which ranks sentences by lexical relevance, and a Dense Retriever [13] using semantic embeddings for similarity search. For MER, we resorted to BERTimbau-large [21] for Portuguese, and XLM-RoBERTa-large [10] for English, with and without a CRF layer. Training was conducted for 15 epochs with early stopping (patience of 3 epochs), learning rate 2e-5, batch size 2 per device with gradient accumulation over 4 steps, and weight decay 0.01. We applied a global document-level split: 60% training, 20% validation, and 20% testing, maintaining diversity across municipalities and entity types, and avoiding overlap between minutes. Finally, these approaches were compared with Gemini and Phi utilizing the LangExtract library [9]. More details such as the prompt and hardware used are available in the repository.
5 Results
MBD. XLM-RoBERTa clearly outperforms BM25 and dense retrievers across all metrics and languages (Table 1), with a slight performance drop in English likely due to translation noise and cross-lingual transfer limitations.
Table 1: Results on the QA (Portuguese and English).
Model F1-score EM
pt en pt en
BM25 0.094 0.093 0.0 0.021
Dense Retriever 0.077 0.111 0.0 0.083
XLM-Roberta 0.826 0.714 0.792 0.604
MER Table 2 reports the MER results on the global split. All transformer-based models achieve strong performance, with BERTimbau obtaining the best results in Portuguese (0.96) and XLM-RoBERTa+Deslex in English (0.94). Entity-level results are not tabulated, but are available in our repository. Overall, the models perform particularly well on core metadata categories such as date, start/end time, and meeting number, which are extracted with near-perfect accuracy (F1 = 1.00). A detailed error analysis shows that most remaining errors are boundary-related, indicating that the models generally identify the correct entity type but struggle with precise span delimitation. In contrast, participant-related entities exhibit higher variability: while “councilors present” reaches a F1-score of 0.98, absent and replaced members range between 0.28 and 0.70, with frequent confusion between presence states (present, absent, and substituted). Meeting type categories are similarly imbalanced, with “ordinary meetings” being robustly detected (F1 = 0.95–1.0), whereas “extraordinary meetings” underperform due to the very limited number of examples (only six instances).
Table 2: Results on the global split (Portuguese and English).
Model F1-score Precision Recall
pt en pt en pt en
BERTimbau 0.96 0.94 0.96 0.94 0.96 0.94
XLM-RoBERTa 0.93 0.94 0.94 0.94 0.92 0.93
BERTimbau + CRF 0.95 0.93 0.95 0.92 0.96 0.95
XLM-RoBERTa + CRF 0.90 0.89 0.88 0.88 0.93 0.91
BERTimbau + Deslex 0.96 0.93 0.96 0.93 0.96 0.94
XLM-RoBERTa + Deslex 0.96 0.94 0.96 0.94 0.97 0.94
Leave-one-out results (Table 3) show reduced generalization compared to the global split (F1 = 0.80 in Portuguese and F1 = 0.72 in English). Detailed error analysis reveals that this degradation is primarily driven by an increase in boundary errors, which become more frequent with deslexicalization in both languages. In Portuguese, the performance drop is mainly reflected in lower recall, indicating missed entities after the removal of lexical cues, whereas in English, the decrease is largely associated with overgeneralization and less precise span boundaries. In incremental evaluation (Table 4), performance improves sharply after the first annotated minutes, with most municipalities reaching F1-scores above 0.95 by the third iteration, showing that minimal local supervision enables rapid model specialization without full retraining.
Table 3: Leave-one-out results for the Portuguese (left) and English (right) models tested without (Base) and with (Deslex) deslexicalization.
BERTimbau F1 Precision Recall
Base 0.80 0.79 0.81
Deslex 0.78 0.79 0.79
XML-Roberta F1 Precision Recall
Base 0.72 0.73 0.72
Deslex 0.70 0.70 0.72
Table 4: Incremental evaluation results per municipality (F1-score).
Examples Alandroal Campo Maior Covilhã Fundão Guimarães Porto
0 0.662 0.758 0.802 0.899 0.784 0.820
1 0.960 0.995 0.872 0.980 0.915 0.971
2 0.977 0.995 0.923 0.986 0.917 0.999
3 0.995 0.995 0.929 0.989 0.941 0.989
4 0.995 0.995 0.953 0.989 0.949 0.993
5 0.995 0.995 0.947 0.993 0.963 0.996
We also compare our pipeline with general-purpose LLMs. While LLM comparison is included, our primary goal is to evaluate the MiNER pipeline rather to exhaustively benchmark LLMs. Thus, we focus on open-weight (Phi) and freely accessible (Gemini) models to ensure reproducibility and enable practitioners with limited budgets to replicate our findings. The Gemini model achieved a modest F1-score (0.27), characterized by high precision (0.83) but low recall (0.16), indicating that, while it correctly identifies some entities, it often fails to capture their structure in municipal minutes, leading to fragmented outputs. The Phi model could not be reliably evaluated because of inconsistent JSON outputs, though structured decoding techniques could mitigate this limitation. In contrast, the pipeline (MBD + NER) consistently produced fully structured outputs, achieving substantially higher performance while requiring over 1,800 times less inference time (approximately 0.4 s vs. 737 s) and emitting nearly 400 times lower carbon emissions ($6\times 10^{-6}$ vs. $2.2\times 10^{-3}$ kg CO2e). These results highlight that, despite their flexibility, the evaluated LLMs are less suited than task-specific models for structured metadata extraction in this domain.
Finally, we conducted an ablation experiment to assess the impact of the first stage of the pipeline. When the NER component was applied directly to the full documents, without the first stage, the F1-score slightly decreased from 0.965 to 0.945. Although this difference is relatively small, the MBD stage plays a crucial role in practice by substantially reducing the amount of text processed by the NER model, decreasing the training time from approximately 50 minutes to only 1 minute. Moreover, by restricting NER to semantically relevant regions in such documents, the segmentation step reduces noise, and enables the reuse of the extracted segments for other downstream tasks that rely on accurate document boundaries, such as meeting summarization, text segmentation, or voting identification.
6 Conclusion and Future Work
In this paper, we present MiNER, a two-stage pipeline for metadata extraction from municipal meeting minutes, which combines QA-based boundary detection with transformer-based NER models. MiNER achieves strong results across Portuguese and English datasets, demonstrating that domain-specific fine-tuning can outperform large general-purpose LLMs while remaining computationally efficient. Despite its strengths, MiNER’s generalization across municipalities remains limited, as stylistic and structural variability hinders transfer learning. Moreover, the segmentation stage, although beneficial for noise reduction, offers only marginal gains in end-to-end accuracy. Future work will explore adaptive fine-tuning strategies, domain-agnostic representations, and integration with summarization or document-linking pipelines to extend MiNER’s applicability to broader information management tasks.
{credits}
Preprint and Version of Record
This preprint has not undergone peer review (when applicable) or any post-submission improvements or corrections. The Version of Record of this contribution is published in Advances in Information Retrieval. ECIR 2026. Lecture Notes in Computer Science, vol 16484. Springer, Cham, and is available online at https://doi.org/10.1007/978-3-032-21300-6_33.
Acknowledgements
This work is financed by National Funds through the FCT - Fundação para a Ciência e a Tecnologia, I.P. (Portuguese Foundation for Science and Technology) within the project CitiLink, with reference 2024.07509.IACDC/2024 (DOI https://doi.org/10.54499/2024.07509.IACDC) within the scope of PRR funding, investment “RE-C05-i08 - More Digital Science”.